Angle A has the property that twenty-six times its compliment is equal to eleven times its supplement. Find the measure of angle A.
Answer by htmentor(789)

The complement of angle A = 90 - A
The supplement of angle A = 180 - A
Given: 26 times the complement = 11 times the supplement:
26(90-A) = 11(180-A)
(26/11)*90 - (26/11)A = 180 - A
(26/11 - 1)A = 90(26/11 - 2)
15/11A = (4/11)90
A = (4/15)90 = 4*6 = 24